OpenStudy (anonymous):

there is a circular path around a sports field.geeta takes 20 minutes to drive one round of the field.while ravi takes 14 minutes for the same.suppose they both start at the same point and at the same direction .after how many minutes will they meet again at starting point

OpenStudy (anonymous):

It'll be the least common multiple of those two times.

OpenStudy (anonymous):

ans is??????????

OpenStudy (anonymous):

Did you even try the problem?

OpenStudy (anonymous):

ya but its not coming

OpenStudy (anonymous):

What is the least common multiple of 14 and 20?

OpenStudy (anonymous):

the ans is 42 but im not getting

OpenStudy (anonymous):

Its finding the Least Common Multiple of 20 and 14. 20= 2x2x5 and 14=2x7. so 2x2x5x7 = 140

OpenStudy (anonymous):

The LCM is not 42... look at the prime factorizations:\[14=2\cdot 7\]\[20 = 2^2 \cdot 5\]

OpenStudy (anonymous):

then wat is the correct ans?

OpenStudy (anonymous):

You're really not attempting this are you...

OpenStudy (anonymous):

oh im getting 140? rit?

ganeshie8 (ganeshie8):

correct.

OpenStudy (anonymous):

thnx